# What is my quantum computer good for? Quantum capability learning with physics-aware neural networks

### v1

@@ 4.3 Encoding schemes
We used two different encoding schemes for converting each circuit $c$ into a tensor. For the CNNs on experimental data, we used the same encoding scheme as Hothem et al. [2023c], as we used their data and networks. For all qpa-NNs, and the CNNs on simulated data, we used the following scheme. As outlined in Section 3, each width-$w$ circuit $c$ is represented by a three-dimensional tensor $I(c)\in\{0,1\}^{n\times d(c)\times n_{ch}}$ describing the gates in $c$ and a matrix $M(C)\in\{0,1\}^{2\times w}$ describing the measurement of the qubits. The $ij$-th entry of $I(c)$,
@@
5 Experiments
We now present the results from our head-to-head comparison between the qpa-NNs and the CNNs on the 5-qubit datasets used in Hothem et al. [2023c]. Figure 2 shows the mean absolute error (MAE) achieved by the CNNs (${\color[rgb]{1,.5,0}\bm{+}}$) and the qpa-NNs (${\color[rgb]{0,0,1}\bm{\bullet}}$) on each of the datasets. For all datasets, MAE is lower for the qpa-NNs than the CNN, with an average reduction of $50.4\%$ ($\sigma_{x}=16.7\%$, i.e., the standard deviation of percent-drop in MAE). The Bayes factor $K$ is between $K=10^{30}$ and $K=10^{383}$ (here, $K$ is the ratio of the likelihood of the qpa-NN to the likelihood of the CNN given the test data). This is overwhelming evidence that the qpa-NN is a better model ($K\geq 10^{2}$ is typically considered decisive). These results strongly suggest that the extra infrastructure in the qpa-NNs is making a difference.
The improved performance of the qpa-NNs is not because of an increase in model size. For example, the $\mathtt{ibmq\_london}$ CNN contains $6,649,531$ trainable parameters compared to the $1,218,348$ trainable parameters in the qpa-NN. Moreover, CNNs of similar or larger sizes than the qpa-NNs were included in the hyperparameter optimization space of the CNNs [Hothem et al., 2023c].
@@ 5 Experiments
6 Simulations
Figure 3: Demonstrating our qpa-NNs’ accuracy for hard-to-model coherent errors. (a) Scatter plot of the prediction errors on test data of a qpa-NN and CNN trained to predict the fidelity $F(c)$ of random circuits run on a hypothetical 4-qubit quantum computer. The qpa-NN significantly outperforms the CNN. The top subplot contains a histogram (green bars) of the ground-truth fidelities. (b) Prediction errors on out-of-distribution test data, from random mirror circuits. The qpa-NN achieves modest prediction accuracy on this out-of-distribution task, suggesting that the qpa-NNs are accurately learning error rates.
One reason why the extra infrastructure in our qpa-NNs may be necessary is that off-the-shelf networks struggle with modeling coherent errors [Hothem et al., 2023c]. To test our hypothesis, we trained a qpa-NN to predict the fidelity $F(c)$ of random circuits executed on a hypothetical 4-qubit quantum computer experiencing purely coherent errors. We compared this qpa-NN to a hyperparameter-tuned CNN trained on the same data. Figure 3 shows the results from one representative dataset.
@@ Appendix A Compute resources
All of the quantum-physics-aware neural networks were trained using a 6-Core Intel Core i9 processor on a MacBookPro 15.1 with 32GB of memory. Each model took roughly 15-20 wall clock minutes to train. Total training time, across the paper, totaled $\sim 160$ wall clock minutes.
All of the simulations and data pre-processing were performed using a 6-core Intel Core i9 processor on a MacBookPro 15.1 with 32GB of memory. Each dataset took approximately 1 hour of wall clock time to create. This total includes the initial circuit creation, simulating the circuits, and encoding each circuit into a tensor.
@@ Appendix B Code and data availability
The simulated 5-qubit data as well as records of all the quantum physics-aware networks can be found in Anonymous . The CNNs and 5-qubit experimental datasets used in Hothem et al. [2023c] are available at Hothem et al. [2023a]. The datasets were originally located at Proctor et al. [2021b]. Each dataset was released under a CC-BY 4.0 International license.
@@ Appendix C Datasets
Table 1: Summary data of every dataset used in the paper. The data for the ring processors is averaged over the 5 simulated datasets. See Figure 4 for images of each processor geometry (i.e., the qubit connectivity graph).
We provide additional details on the datasets used in the paper. Table 1 summarizes each dataset. We tracked all weight-$2$ errors with support on qubits connected by $2$ hops in all the datasets. Below, we provide additional details on the circuit and error model generating processes.
@@ C.1 Creating the circuits
Figure 4: Device geometries. The connectivity graphs for the (a) 5-qubit $\mathtt{ibmq\_yorktown}$ “bowtie” processor; (b) the remaining 5-qubit experimental “t-bar” processors; and (c) the 4-qubit simulated “ring” processor.
@@
C.2 Creating an error model
@@ D.2 Convolutional neural network details
Details on the specific convolutional neural networks used in this paper are located in Hothem et al. [2023c]. We fine-tuned each network on high-PST experimental data using the Adam optimizer and early stopping.
@@ Appendix F Simulation results
Table 4: Summary model performance data on simulated data.
@@
Table 4 summarizes model performance on each of the simulated datasets used in the paper. Copies of the pre-trained quantum-physics-aware neural networks and fine-tuned CNNs are available in the Supplementary Material. Scatter plots of each model’s prediction errors are also available in the Supplementary Material.

### v2

@@ 4.4 Encoding schemes
We used two different encoding schemes for converting each circuit $c$ into a tensor. For the CNNs on experimental data, we used the same encoding scheme as Hothem et al. [2024b], as we used their data and networks. For all qpa-NNs, and the CNNs on simulated data, we used the following scheme. As outlined in Section 3, each width-$w$ circuit $c$ is represented by a three-dimensional tensor $I(c)\in\{0,1\}^{n\times d(c)\times n_{ch}}$ describing the gates in $c$ and a matrix $M(C)\in\{0,1\}^{2\times w}$ describing the measurement of the qubits. The $ij$-th entry of $I(c)$,
@@
5 5-qubit experiments
We now present the results from our head-to-head comparison between the qpa-NNs and the CNNs on the 5-qubit datasets used in Hothem et al. [2024b]. Figure 2 shows the mean absolute error (MAE) achieved by the CNNs (${\color[rgb]{1,.5,0}\bm{+}}$) and the qpa-NNs (${\color[rgb]{0,0,1}\bm{\bullet}}$) on each of the datasets. For all datasets, MAE is lower for the qpa-NNs than the CNNs, with an average reduction of $50.4\%$ ($\sigma_{x}=16.7\%$, i.e., the standard deviation of percent-drop in MAE). The Bayes factor $K$ is between $K=10^{30}$ and $K=10^{383}$ (here, $K$ is the ratio of the likelihood of the qpa-NN to the likelihood of the CNN given the test data). This is overwhelming evidence that the qpa-NN is a better model ($K\geq 10^{2}$ is typically considered decisive). These results strongly suggest that the extra infrastructure in the qpa-NNs is making a difference.
The improved performance of the qpa-NNs is not because of an increase in model size. For example, the $\mathtt{ibmq\_london}$ CNN contains $6,649,531$ trainable parameters compared to the $1,218,348$ trainable parameters in the qpa-NN. Moreover, CNNs of similar or larger sizes than the qpa-NNs were included in the hyperparameter optimization space of the CNNs [Hothem et al., 2024b].
@@
6 4-qubit simulations
Figure 3: Demonstrating our qpa-NNs’ accuracy for hard-to-model coherent errors and at scale. (a) Scatter plot of the prediction errors on test data of a qpa-NN (${\color[rgb]{0,0,1}\bm{\bullet}}$) and CNN (${\color[rgb]{1,.5,0}\bm{\bullet}}$) trained to predict the fidelity $F(c)$ of random circuits run on a hypothetical 4-qubit quantum computer. The qpa-NN significantly outperforms the CNN. The top subplot contains a histogram (green bars) of the ground-truth fidelities. (b) Prediction errors on out-of-distribution test data, from random mirror circuits. The qpa-NN achieves modest prediction accuracy on this out-of-distribution task, suggesting that the qpa-NNs are accurately learning error rates. (c) Prediction errors on the 100-qubit test data, demonstrating that our qpa-NN approach can accurately predict $F(c)$ for circuits run on large-scale quantum computers.
One reason why the extra infrastructure in our qpa-NNs may be necessary is that off-the-shelf networks struggle with modeling coherent errors [Hothem et al., 2024b]. To test our hypothesis, we trained a qpa-NN to predict the fidelity $F(c)$ of random circuits executed on a hypothetical 4-qubit quantum computer experiencing purely coherent errors. We compared this qpa-NN to a hyperparameter-tuned CNN trained on the same data. Figure 3 shows the results from one representative dataset.
@@
7 100-qubit simulation
Quantum-physics-aware neural networks scale just as well as CNNs, despite their extra infrastructure. To demonstrate their scalability, we trained a qpa-NN to predict the fidelity $F(c)$ of random circuits executed on a hypothetical 100-qubit quantum computer experiencing a mix of stochastic and coherent errors. To our knowledge, this is the first creation of a capability model of any kind, for a 100+ qubit quantum computer. Figure 3(c) shows the results from our demonstration.
The qpa-NN achieved a MAE of $0.097\%$. While the underlying noise model was quite simple, this result shows that it is technically feasible to construct qpa-NN capability models for today’s moderate-scale quantum computers and for tomorrow’s early fault-tolerant quantum computers.
@@ Appendix A Compute resources
All of the quantum-physics-aware neural networks used in our 5-qubit experiments and 4-qubit simulations were trained using a 6-Core Intel Core i9 processor on a MacBookPro 15.1 with 32GB of memory. Each model took roughly 15-20 wall clock minutes to train. Total training time, across the paper, totaled $\sim 160$ wall clock minutes.
All of the 4-qubit simulations and data pre-processing were performed using a 6-core Intel Core i9 processor on a MacBookPro 15.1 with 32GB of memory. Each dataset took approximately 1 hour of wall clock time to create. This total includes the initial circuit creation, simulating the circuits, and encoding each circuit into a tensor.
All of the 100-qubit simulations, data pre-processing, and model training were performed using two 14-core Intel Xeon CPU E5-2697 v3 @ 2.60GHz processors. An end-to-end run (i.e., circuit generation to trained model predictions) took roughly 12 hours of wall clock time.
@@ Appendix B Code and data availability
The simulated data as well as records of all the quantum physics-aware networks will be released publicly once they clear Sandia’s copyright process. Until then, please email the authors. The CNNs and 5-qubit experimental datasets used in Hothem et al. [2024b] are available at Hothem et al. [2023a]. The datasets were originally located at Proctor et al. [2021b]. Each dataset was released under a CC-BY 4.0 International license.
@@ Appendix C Datasets
$100$-qubit Ring ring random i.i.d. 100 qubits 1-22 layers 2812 circuits 938 circuits 1250 circuits
Table 1: Summary data of every dataset used in the paper. The data for the 4-qubit ring processors is averaged over the 5 simulated datasets. See Figure 4 for images of each processor geometry (i.e., the qubit connectivity graph).
We provide additional details on the datasets used in the paper. Table 1 summarizes each dataset. We tracked all weight-$2$ errors with support on qubits connected by $2$ hops in all, but the $100$-qubit datasets. Below, we provide additional details on the circuit and error model generating processes.
@@ C.1 Creating the circuits
Figure 4: Device geometries. The connectivity graphs for the (a) 5-qubit $\mathtt{ibmq\_yorktown}$ “bowtie” processor; (b) the remaining 5-qubit experimental “t-bar” processors; and (c) the 4-qubit simulated “ring” processor. The 100-qubit simulated “ring” processor has the same topology, just more qubits.
@@
C.2 Creating a 4-qubit error model
@@
C.3 Creating a 100-qubit error model
In this subsection, we explain how we constructed the 100-qubit Markovian error model used in Section 7. As with the 4-qubit error model, we provide a conceptual explanation. The actual error model was created in $\mathtt{pygsti}$.
As with the 4-qubit error model, the 100-qubit Markovian error model was specified using the error generator framework explained in Section 2. The error model consisted of gate-dependent errors sampled according to a two-step process. Unlike with the 4-qubit error model, the 100-qubit error model included non-local Pauli stochastic and coherent errors, and the errors for each gate were independent of the qubit acted upon by the gate. Moreover, all of the errors in the 100-qubit error model are weight-1 errors (i.e., they affect a single qubit).
Each gate’s error strengths were independently sampled. First, we enumerated all 600 possible weight-1 Pauli stochastic and coherent errors in a 100-qubit device. Then, for each gate $g$, we independently sampled the strength of each of the 300 weight-1 Pauli stochastic errors and the 300 possible weight-1 coherent errors. The strengths were sampled uniformly random, with a maximum Pauli stochastic error strength of $0.0000001$ and a maximum coherent error strength of $0.00005$. The resulting 600 error strengths were assembled into a 600-dimensional error vector for the gate, $\vec{\varepsilon}_{g}$.
C.4 Simulating the 100-qubit circuits
In this subsection, we describe the first-order simulation method used to approximate $F(c)$ in our 100-qubit simulations (Section 4.3). The method works by constructing an approximate error matrix $E(c)$ [Fig. 1(e)], and then estimating $F(c)$ by performing the same computation as in the second half of a qpa-NN [Fig. 1(f)]. Herein, we describe how to construct the approximate error matrix $E(c)$. Readers should refer to Section 3 for an in-depth explanation on how to use $E(c)$ to estimate $F(c)$.
For each circuit $c$, we constructed an approximate error matrix $E(c)$ by concatenating approximate error vectors $E_{i}(c)$ for each circuit layer in $c$. Each layer’s error vector was computed as a linear combination of the individual gate error vectors, with coefficients equal to the number of times each gate appears in the circuit layer.
@@ D.2 Convolutional neural network details
Details on the specific convolutional neural networks used in this paper are located in Hothem et al. [2024b]. We fine-tuned each network on high-PST experimental data using the Adam optimizer and early stopping.
@@ Appendix F Simulation results
Table 4: Summary model performance data on the 4-qubit simulated data.
@@
Table 4 summarizes model performance on each of the 4-qubit simulated datasets used in the paper. Copies of the pre-trained quantum-physics-aware neural networks and fine-tuned CNNs are available in the Supplementary Material. Scatter plots of each model’s prediction errors are also available in the Supplementary Material.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, we list the main claims and the section in which they are answered below.
    1. (a)
      We developed a new physics-aware neural network architecture for quantum capability learning: Section 3.
    2. (b)
      Our new approach achieves up to a $\sim 50\%$ and $\sim 76\%$ improvement over state-of-the-art convolutional neural networks on experimental and simulated data, respectively: Section 5 and 6.
    3. (c)
      Our new approach beats state-of-the-art convolutional networks, in part, due to their improved ability to model coherent errors: Section 6.
    4. (d)
      Our new approach achieves moderate prediction accuracy on an out-of-distribution prediction task: Section 6.
    5. (e)
      Our approach scales to 100+ qubits.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We include a discussion of the limitations of the work in Section 8.1.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not include any new theoretical results or proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We believe that we provide sufficient details to reproduce the main experimental results of the paper. Readers should be able to recreate our results based on the details in the main body of the paper and the appendix, or by using the notebooks in the supplemental material once released.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We plan to provide open access to all the data and code used in the paper through a GitHub repository. We also provide explicit instructions on how to access the experimental data in Appendix B.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Section 4 we explain how we processed the experimental and simulated datasets. Appendices C and D contain additional details on the datasets, specific network instantiations, and model training.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report error bars on the mean absolute error of the models’ predictions in Section 6. We do not report error bars on the models’ predictions in the experimental data as the original paper [Hothem et al., 2024b] reported their error bars as being trivial. However, we do provide the standard deviation of the percent change in the MAE across the experimental datasets, and we report Bayes factors for each model on each experimental dataset, demonstrating substantial improvement by the physics-aware networks.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide details on the compute resources and compute time used in this work in Appendix A.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We believe that we have conducted our research in a manner that conforms, in every respect, with the NeurIPS Code of Ethics. The only relevant areas of concern are the use of deprecated datasets and respect for copyright and fair use. We believe that we have not violated either of these requirements, although the cloud-accessed processors used in the experimental section are no longer available (but the datasets are not deprecated).
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: We believe that there is little to no societal impact of our work. We believe that the following quote from the checklist guidelines is relevant: “it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.” While quantum computers might some day have a large societal impact, our work does not directly improve their ability to run programs with societal impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We explicitly reference each existing dataset in the References, and state each existing datasets license in Appendix B. All existing assets were licensed under a CC-BY 4.0 license, requiring proper attribution.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: At the time of submission, we have released code used in Section 6 and the new simulated dataset in the supplemental material. In the future we will release additional assets, such as the models trained on experimental data, publicly after receiving approval from our employer.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.